OpenStudy (anonymous):

Taylor polynomial??? Lagrange remainder?? I dont know how to find Taylor polynomial P4 of f(x)= x + 3sin(x) and Lagrange form of the remainder R3(x) of f(x)= ln(3x^2) + 6x^4 Please please i'm so lost in this part of calculus

OpenStudy (anonymous):

Hi nikets, for first question about taylor polonomial, my answer is sinx = x/1! - x^3/3! + x^5/5!.. 3sinx =3*(x/1! - x^3/3! + x^5/5!...) x + 3sinx = 4x -x^3/2 + x^5/5*4*2 + ...

OpenStudy (anonymous):

4x-x^3/2 + o[x]^4